OpenStudy (anonymous):

Lynda has a standard deck of 52 playing cards. Shuffles the deck, then draws a card without looking. a) What is the probability Lynda draws a red card? b) What is the probability Lynda draws a 4? c) What is the probability Lynda draws the 4 of diamonds?

OpenStudy (anonymous):

....Two Lynda? I'm just going to ignore the two. (Question number?) There are 26 red cards and 26 blue cards. 26/52=1/2 There are 4 cards of '4' (Spade Diamond Heart Clover) 4/52=1/13 There is one 4 of diamonds in a deck 1/52=1/52